Length of a minute hand of a clock is 4.5cm.find the average velocity of the tip of minute hand between 6 A.M to 6.30 A.M and 6 A.M to 6.30P.M
 

Grade:11

5 Answers

Arun
5 years ago
Between 6 am to 6:30 am
Average speed = Total distance/ Time
= πr / (30 × 60 seconds)
= 4π/1800
= 0.00698 cm/s
Average velocity = Total displacement/ Time
= 2r / (30 × 60 seconds)
= 9 / 1800
= 0.005 cm/s
_____________________________
Average velocity and average speed between 6 am to 6:30 pm is same as that of 6 am to 6:30 am
 
Khimraj
5 years ago
 
Average speed is given by Total distance divided by Time
speedavg = πr /1800
= π*4.5/1800
= 0.00785 cm/s
Average velocity will be given by  Total displacement divided by Time
vavg = 2r / 1800
= 9 / 1800
= 0.005 cm/s
Hope it clears.
Swati
5 years ago

Displacement is the shortest route between 2 points therefore here displacement =2r= 2x4.5

For 6 am to 6:30 am

t=30 min = 30x60

                 =1800s

Therefore, v= 2x4.5

                        1800

              =5x10^-3 cm/s

 

For 6 am to 6:30 pm

t=12.5hr =12x60x60+30x60

               =45000s

Therefore, v=2x4.5

                       45000

                     =2x10^-4 cm/s

Kriti Prakash
4 years ago
Part 1
Vav= displacement ÷ time 
Vav=(4.5+4.5)÷(30×60)= 1÷200 cm/ s
 Part 2
Vav= displacement ÷time
Vav= (4.5+4.5)÷(12hours×60×60 + 30min×60)
Vav = 1÷5000 cm / s
 
Vav= 
Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your question.
 
Displacement is the shortest route between 2 points therefore here displacement =2r= 2x4.5
For 6 am to 6:30 am
t = 30 min = 30x60
                 =1800s
Therefore, v = 2x4.5/1800
                      = 5x10-3 cm/s
 
For 6 am to 6:30 pm
t = 12.5hr = 12x60x60 + 30x60
                  = 45000s
Therefore, v = 2x4.5/45000
                     =2x10-4 cm/s
 
Thanks and regards,
Kushagra